Razor:

The length of a rectangle is 3 more inches than the width. find the length and width of the rectangle if the perimeter of the rectangle is 54 inches. Write an equation that represents the situation. Explain any variable used. Solve the equation. Show your work. State your answer as a complete sentence. I already practically have the answer I just need someone to double check. I won't post my answer's until someone is willing to respond/help and solve the problem with me.

Razor:

>..>

Kaihalojr:

wutchu talkin bout?

Kaihalojr:

help dis boi e.e

Razor:

help meh

kittybasil:

wot

kittybasil:

"The length of a rectangle is 3 more inches than the width..." if \(l\) is length and \(w\) is width;\[l=3+w\] "...the perimeter of the rectangle is 54 inches" perimeter is \(2l+2w\), so\[2l+2w=54\] now we translate together by plugging in the length expression for \(l\) (substitution):\[2(3+w)+2w=54\]and simplify...\[6+2w+2w=54\]\[4w=54-6\]\[4w=48\]\[w=48\div4=12\]your width is \(\color{red}{12}\) inches. to find length: add 3 to the width, like the equation...\[l=3+\color{red}{12}=15\] 15 inches.

Razor:

.......................................THANK YEWWWWWWWWWWWW

Razor:

LOL

Kaihalojr:

Kaihalojr:

pika lenny for the road

kittybasil:

ಠ_ಠ ok good luck on your homework razor.

Kaihalojr:

kittybasil:

emoji wars in the chat only pls kthxbye